# On the analog category of finite groups

### v1

@@ Abstract.
We show that, up to small error, the analog category of a finite group records the size of its largest Sylow subgroup.
@@ 1. Introduction
If $G$ is finite, then $\mathrm{acat}(G)<|G|$.
We show here that this bound is arbitrarily far from sharp and almost never holds. Provisionally, let us call $G$ $a$-special if $|G|=ap^{s}$ with $p$ a prime, $(a,p)=1$, and $p^{s}>a$ (thus, a $1$-special group is simply a $p$-group).
@@
Let $G$ be a finite group.
@@
    If $G$ is not $a$-special for some $a\in\{1,2,3\}$, then the universal upper bound does not hold for $G$.
@@
    If $G$ is $1$- or $2$-special, then the universal upper bound holds for $G$ [^1] Unfortunately, we do not know whether the universal upper bound holds for $3$-special groups.
@@
    For any $N\geq 0$, there is a group $G$ such that $\mathrm{acat}(G)<|G|-N$.
In other words, the universal upper bound holds for $p$-groups and fails for almost all groups not of prime power order. For groups of the latter type, we show that, up to small error, the analog category is recording the size of the largest Sylow subgroup.
@@
$\min\left\{2,\textstyle{\frac{|N(P)|}{|P|}}\right\}\leq\frac{\mathrm{acat}(G)+1}{|P|}\leq 3.$
@@
Prior to our work here, the quantity $\mathrm{acat}(BG)$ was almost completely unknown apart from the universal upper bound and a calculation for cyclic groups of prime order [Dra]. Strictly speaking, this last calculation was of an a priori different invariant, the distributional category; we show here that the two coincide for finite groups, a special case of [KW24, Conj. 1.2].
@@ 1.1. Conventions
We use the following non-standard notational convention for topological simplices:
@@
We work in Steenrod’s convenient category of topological spaces [Ste67]—see [KW24, Appendix A] for a summary of relevant facts about these spaces. Topological spaces are implicitly convenient, as are limits, including products, and mapping spaces. Convenient colimits, when they exist, are the same as ordinary colimits. The adjective “compact” refers to the definition in terms of open covers. We write $BG$ for the classifying space of the (discrete) group $G$, i.e., the geometric realization of its nerve. We write $EG$ for the universal cover of $BG$ and $X^{hG}=\mathrm{Map}^{G}(EG,X)$ for the space of homotopy fixed points of the $G$-space $X$.
@@ 2. The analog category of a group
The purpose of this section is to establish the following formula, which the reader may take as a definition of the analog category $\mathrm{acat}(G)$.
@@
[…]
@@
Since $X$ is contractible, an equivariant map $EG\to X$ exists, which is to say that $X^{hG}\neq\varnothing$. It follows that $(\Delta^{G}_{n})^{hG}$ receives a map from a non-empty space, hence is itself non-empty. The claim follows from Proposition 2.1. ∎
@@ 3. Designer complexes
We proceed by simultaneous induction on $d$ and $d_{p}(G)$. In the case $d=d_{p}(G)=0$, we let $X_{p}(G)_{0}$ be any connected $1$-dimensional free $G$-complex. Notice that the third condition is vacuous in this case. For $d=0$ and general $G$, we choose a $p$-Sylow subgroup $P$ and set
@@
The second property holds by construction, and the third follows from the observation that the fixed set of any $p$-Sylow subgroup is homeomorphic to $\overline{X}_{p}(N(P)/P)_{0}$, which is $p$-acyclic by construction.
@@ 4. Proofs of the main results
$\left(\Delta^{G}_{n}\right)^{H}\cong\Delta^{G/H}_{\lfloor\frac{n+1}{|H|}\rfloor-1}$
@@
As the restriction of a linear map, this function is continuous, and its image is $H$-fixed by inspection; thus, we may we view $f$ as a map to $(\Delta^{G})^{H}$. As such, it is $N(H)/H$-equivariant and injective by inspection, and we claim that it is also surjective. To see why, note that a point in $\Delta^{G}$ is fixed by $H$ if and only if the barycentric coordinate of $gh$ is independent of $h\in H$ for every $g\in G$. Thus, given $t\in(\Delta^{G})^{H}$, the assignment $gH\mapsto|H|t_{g}$ is a well-defined element of $f^{-1}(t)$. The homeomorphism $(\Delta^{G})^{H}\cong\Delta^{G/H}$ follows, since both sides are compact and Hausdorff.
@@
In the following, we write $X_{p}^{\wedge}$ for the completion of the space $X$ at the prime $p$—see [MP11], for example. We recall that, according to the Sullivan conjecture, the natural map $(X^{P})_{p}^{\wedge}\to(X_{p}^{\wedge})^{hP}$ is a weak equivalence for any $p$-group $P$ and finite $P$-CW complex $X$.
@@
For the first claim, it suffices to show that $(\Delta_{n}^{G})^{hP}=\varnothing$, since $(\Delta_{n}^{G})^{hG}$ has a canonical map to this space. By the Sullivan conjecture, the second of the following canonical maps is a weak equivalence:
$(\Delta_{n}^{G})^{P}\to((\Delta_{n}^{G})^{P})^{\wedge}_{p}\to((\Delta_{n}^{G})^{\wedge}_{p})^{hP}\leftarrow(\Delta_{n}^{G})^{hP}.$
If $n<|P|-1$, then $(\Delta^{G}_{n})^{P}=\varnothing$ by Proposition 4.1. In particular, this space is $p$-complete, so the first map above is also a weak equivalence. We conclude that the target of the rightmost map is empty, so its source must be so as well.
If $|P|-1\leq n<2|P|-1$, then Proposition 4.1 instead identifies $(\Delta_{n}^{G})^{P}$ with the discrete $N(P)/P$-space $G/P$, which is also $p$-complete. We consider the result of applying homotopy fixed points for $N(P)/P$ to the above string of arrows. Since $P$ is not self-normalizing, we again obtain the empty space in the leftmost position, implying this time that $(\Delta_{n}^{G})^{hN(P)}=((\Delta_{n}^{G})^{hP})^{hN(P)/P}=\varnothing$. Since $(\Delta_{n}^{G})^{hG}$ maps to this space, the conclusion follows. ∎
@@
If $r\notin\{1,2,3\}$, then the righthand side of this inequality is strictly less than $1$, implying the first claim. The second claim is immediate from Proposition 4.3 and the universal upper bound. For the third claim, fix $N>0$ and a prime $p>5$, and take $s$ sufficiently large so that $N\leq 2p^{s}$. In this case, we have $\mathrm{acat}(C_{5p^{s}})<3p^{s}=|G|-2p^{s}\leq|G|-N.$
@@
The goal of this section is to prove Theorem 1.3, claiming that the analog category of the finite group $G$ coincides with its distributional category in the sense of [DJ24]. We begin by recalling the relevant definitions. [^3] For the sake of easier reading, we depart from the notation of [DJ24].
@@
Strictly speaking, the definition of distributional sectional category is not contained in [DJ24], but it should be regarded as implicit. An issue requiring slightly more care is that the distributional category is not defined if $X$ is not metrizable—for example, $X=BG$ with $G$ infinite. One possible workaround is to appeal to the fact that any CW complex is metrizable up to homotopy equivalence [Cau74]. Fortunately, since we confine our discussion here to finite groups, the issue does not arise.
@@
Note that the first claim of Lemma 5.3 is simply the claim that the quotient topology on $\mathcal{P}(X)$ is finer than the Lévy–Prokhorov topology when both are defined.
@@
which is to say sequential continuity, since the source is metric. Since $X$ is separable, because compact, the topology on the target is the topology of weak convergence of measures, so sequential continuity follows from continuity of the composite
@@
[…]

### v2

@@ Abstract.
We show that the analog category of a finite group is essentially proportional to the size of its largest Sylow subgroup. We conclude that the universal upper bound given by the order of the group is very far from optimal.
@@ 1. Introduction
If $G$ is finite, then $\mathrm{acat}(G)+1\leq|G|$. [^1] For stating our results, it is convenient to work with the “unreduced” category, which differs from the “reduced” convention by $1$. To avoid confusion, we will simply write $\mathrm{acat}(G)$ for the latter and $\mathrm{acat}(G)+1$ for the former. Both conventions are common throughout the literature on Lusternik–Schnirelmann category and topological complexity.
We show here that this bound is a very bad one in three senses: the groups for which it is sharp are highly constrained; the difference of the two sides is arbitrarily large; and their ratio is arbitrarily small. Provisionally, let us call $G$ $a$-special if $|G|=ap^{s}$ with $p$ a prime, $(a,p)=1$, and $p^{s}>a$ (thus, a $1$-special group is simply a $p$-group).
@@
In what follows, $G$ refers to a finite group.
@@
    If $G$ is not $a$-special for some $a\in\{1,2,3\}$, then the universal upper bound is strict for $G$.
@@
    If $G$ is $1$- or $2$-special, then the universal upper bound is sharp for GG. [^2] Unfortunately, we do not know whether the universal upper bound is sharp for $3$-special groups.
@@
    For any $N\geq 0$ and $\epsilon>0$, there exists a $G$ such that $\mathrm{acat}(G)+1\leq|G|-N$ and $\mathrm{acat}(G)+1\leq\epsilon|G|$.
In particular, the universal upper bound is sharp for $p$-groups and strict for almost all other groups. For groups of the latter type, we show that the analog category is roughly proportional to the size of the largest Sylow subgroup.
@@
$\min\left\{2,\textstyle{\frac{|N(P)|}{|P|}}\right\}\leq\frac{\mathrm{acat}(G)+1}{|P|}\leq 3,$
where $N(P)$ denotes the normalizer of $P$ in $G$.
@@
Prior to our work here, the quantity $\mathrm{acat}(G)$ was almost completely unknown apart from the universal upper bound and a calculation for cyclic groups of prime order [Dra]. Strictly speaking, this last calculation was of an a priori different invariant, the distributional category; we show here that the two coincide for finite groups, a special case of [KW24, Conj. 1.2].
@@ 1.1. Conventions
We write $\Delta^{S}$ for the simplex spanned by the set $S$. Thus, a point in $\Delta^{S}$ is a formal sum $\sum_{s\in S}t_{s}s$, where the barycentric coordinates $t_{s}$ are non-negative numbers summing to $1$, all but finitely many of which vanish. In the case $S=\{1,\ldots,n\}$, we make the abbreviation $\Delta^{\{1,\ldots,n\}}=\Delta^{n-1}$; explicitly, this choice amounts to the following slightly non-standard notational convention:
@@
We work in Steenrod’s convenient category of topological spaces [Ste67]—see [KW24, Appendix A] for a summary of relevant facts about these spaces. Topological spaces are implicitly convenient, as are limits, including products, and mapping spaces. Convenient colimits, when they exist, are the same as ordinary colimits. The adjective “compact” refers to the definition in terms of open covers. We write $BG$ for the classifying space of the (discrete) group $G$, i.e., the geometric realization of its nerve. We write $EG$ for the universal cover of $BG$ and $X^{hG}=\mathrm{Map}^{G}(EG,X)$ for the space of homotopy fixed points of the $G$-space $X$, where $\mathrm{Map}^{G}$ denotes the space of $G$-equivariant maps.
@@ 2. The analog category of a group
The purpose of this section is to establish the following formula, which the reader may take as a definition of the analog category $\mathrm{acat}(G)$. We recall that $\Delta^{G}$ denotes the simplex spanned by the elements of $G$, and $\Delta^{G}_{n}$ is its $n$-skeleton.
@@
${\mathrm{Map}^{G}(EG,EG\times X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Map}^{G}(EG,EG\times_{G}X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\mathrm{Map}(BG,EG\times_{G}X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Map}^{G}(EG,EG)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Map}^{G}(EG,BG)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\mathrm{Map}(BG,BG)}$
@@
Contractibility implies that $X^{hG}\neq\varnothing$, since an equivariant map $EG\to X$ may be constructed by elementary obstruction theory; alternatively, the map $EG\times_{G}X\to BG$ is a trivial fibration, since $EG\to\mathrm{pt}$ is a homotopy equivalence, and Lemma 2.4 applies. It follows that $(\Delta^{G}_{n})^{hG}$ receives a map from a non-empty space, hence is itself non-empty. The claim follows from Proposition 2.1. ∎
@@ 3. Designer complexes
Lemma 3.9.
Let $H\leq G$ be a subgroup and $X$ an $N(H)$-space. For any $g\in G$, there is a canonical homeomorphism
$\left(G\times_{N(H)}X\right)^{gHg^{-1}}\cong X^{H}.$
@@
Consider the standard decomposition $G\times_{N(H)}X=\bigsqcup_{[g_{i}]\in G/N(H)}g_{i}X$, and take $g$ to be one of our coset representatives. Writing $ghg^{-1}g_{i}=g_{j}h^{\prime}$, we have $ghg^{-1}\cdot g_{i}x=g_{j}(h^{\prime}\cdot x)$. It follows that $g_{i}x$ is fixed by $gHg^{-1}$ if and only if $x$ is fixed by $H$ and $g]=[g_{i}]$, which is to say $g=g_{i}$. Thus, the desired homeomorphism is given (from right to left) by $x\mapsto gx$. ∎
Proof of Theorem 3.8.
We proceed by simultaneous induction on $d$ and $d_{p}(G)$. In the case $d=d_{p}(G)=0$, we let $X_{p}(G)_{0}$ be any connected $1$-dimensional free $G$-complex, e.g., a Cayley graph. Notice that the third condition is vacuous in this case. For $d=0$ and general $G$, we choose a $p$-Sylow subgroup $P$ and set
@@
The second property holds by construction, and the third follows from the observation that, by Lemma 3.9, the fixed set of any $p$-Sylow subgroup is homeomorphic to $\overline{X}_{p}(N(P)/P)_{0}$, which is $p$-acyclic by construction.
@@ 4. Proofs of the main results
$\left(\Delta^{G}_{n}\right)^{H}\cong\Delta^{G/H}_{k},$
where $k={\lfloor\frac{n+1}{|H|}\rfloor-1}$.
@@
As the restriction of a linear map, this function is continuous, and its image is $H$-fixed by inspection; thus, we may we view $f$ as a map to $(\Delta^{G})^{H}$. As such, it is $N(H)/H$-equivariant and injective by inspection, and we claim that it is also surjective. To see why, note that a point in $\Delta^{G}$ is fixed by $H$ if and only if the barycentric coordinate of $gh$ is independent of $h\in H$ for every $g\in G$. Thus, given $t\in(\Delta^{G})^{H}$, setting $t_{gH}=|H|t_{g}$ defines an element of $f^{-1}(t)$. The homeomorphism $(\Delta^{G})^{H}\cong\Delta^{G/H}$ follows, since both sides are compact and Hausdorff.
@@
In what follows, we write $X_{p}^{\wedge}$ for the completion of the space $X$ at the prime pp—see [MP11], for example. We recall that, according to one of several results known collectively as the “generalized Sullivan conjecture,” due to Carlsson [Car91, Thm. B(c)] and Dwyer–Miller–Neisendorfer [DMN89], the natural map $(X^{P})_{p}^{\wedge}\to(X_{p}^{\wedge})^{hP}$ is a weak equivalence for any $p$-group $P$ and finite dimensional $P$-CW complex $X$.
@@
Lemma 4.4.
Let $H\leq G$ be any subgroup. If $H$ is not self-normalizing, then $(G/H)^{hN(H)/H}=\varnothing$.
@@
Our assumptions imply that $N(H)/H$ is a nontrivial group acting without fixed point on the discrete space $G/H$. An easy exercise shows that such a space admits no equivariant map from any connected $N(H)/H$-space with nontrivial action, hence no homotopy fixed point. ∎
Proof of Proposition 4.3.
We begin with a few elementary observations regarding the following commutative diagram of canonical maps, to which we will appeal throughout the argument:
${\mathrm{Map}(\mathrm{pt},\Delta^{G}_{n})^{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\left(\mathrm{Map}(\mathrm{pt},\Delta_{n}^{G})^{P}\right)_{p}^{\wedge}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\left(\mathrm{Map}(EG,(\Delta_{n}^{G})_{p}^{\wedge}\right)^{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Map}(EG,\Delta_{n}^{G})^{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\Delta_{n}^{G})^{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${((\Delta_{n}^{G})^{P})^{\wedge}_{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${((\Delta_{n}^{G})^{\wedge}_{p})^{hP}}$${(\Delta_{n}^{G})^{hP}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
First, since the canonical map $EP\to EG$ is a $P$-equivariant homotopy equivalence, the third and fourth vertical arrows are weak equivalences. Second, since $N(P)/P$ acts canonically on the $P$-fixed set of any $G$-space, the arrows in the top row are all equivariant maps between $N(P)/P$-spaces. Third, by the Sullivan conjecture, the second map in the bottom row is a weak equivalence, and hence in the top row as well.
For the first claim of the proposition, it suffices to show that $(\Delta_{n}^{G})^{hP}=\varnothing$ for n<|P|1n<|P|-1, since $(\Delta_{n}^{G})^{hG}$ has a canonical map to this space. In this range, we have $(\Delta^{G}_{n})^{P}=\varnothing$ by Proposition 4.1; in particular, this space is $p$-complete, so the first map in this row is also a weak equivalence (in fact, an equality, but this is irrelevant). We conclude that the target of the rightmost map is empty, so its source must be so as well.
For the second claim, it suffices as before to show that $(\Delta_{n}^{G})^{hN(P)}=\varnothing$ for $|P|-1\leq n<2|P|-1$. In this range, Proposition 4.1 instead identifies $(\Delta_{n}^{G})^{P}$ with the discrete $N(P)/P$-space $G/P$, which is also $p$-complete. As before, it follows that the first arrow in the bottom row of the diagram above is a weak equivalence, and hence in the top row as well. We also conclude from Lemma 4.4 that $(\Delta_{n}^{G})^{P}$ has no homotopy fixed points for the action of $N(P)/P$. Since the homotopy fixed points functor preserves weak equivalences, a diagram chase as in the previous paragraph yields the conclusion that
$\left(\mathrm{Map}(EG,\Delta_{n}^{G})^{P}\right)^{hN(P)/P}=\varnothing.$
An easy exercise shows that this space is weakly equivalent to $(\Delta_{n}^{G})^{hN(P)}$, and the claim follows. ∎
@@
If $r\notin\{1,2,3\}$, then the righthand side of this inequality is strictly less than $1$, implying the first claim. The second claim is immediate from Proposition 4.3 and the universal upper bound. For the third claim, fix $N$ and $\epsilon$ and choose distinct primes $p>q$ and a natural number $s$ so that $N\leq(q-3)p^{s}$ and $\epsilon\leq\frac{3}{q}$. In this case, the desired bounds hold for the cyclic group of order $p^{s}q$. ∎
@@
The goal of this section is to prove Theorem 1.3, claiming that the analog category of the finite group $G$ coincides with its distributional category in the sense of [DJ24, Jau]. We begin by recalling the relevant definitions. [^4] For the sake of easier reading, we depart from the notation of [DJ24].
@@
An issue deserving of care is that the distributional category is not defined if $X$ is not metrizable—for example, $X=BG$ with $G$ infinite. One possible workaround is to appeal to the fact that any CW complex is metrizable up to homotopy equivalence [Cau74]. Fortunately, since we confine our discussion here to finite groups, the issue does not arise.
@@
Note that the first claim of Lemma 5.3 is simply the claim that the quotient topology on $\mathcal{P}(X)$ is finer than the Lévy–Prokhorov topology when both are defined, which can also be seen by considering an explicit basis for the latter [Jau, §8].
@@
which is to say sequential continuity, since the source is metric. Since $X$ is compact, hence separable, the topology on the target is the topology of weak convergence of measures, so sequential continuity follows from continuity of the composite
@@
${\mathcal{P}_{n}(f)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{id}}$${\mathcal{P}_{n}(f)_{\mathrm{LP}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y}$
@@
Remark 5.5.
An alternative argument establishing that the distributional sectional category is a homotopy invariant of fibrations is given in [Jau, Prop. 5.3].